I need to soleve each system of equations by using addition to eliminate one of the variables. Check.
If 3c+d=0
Then c-4d=52

Answer by checkley75(3666)

3c+d=0 multiply this equation by 4 & add.
c-4d=52
12c+4d=0
--------------
13c=52
c=52/13
c=4 ans.
4-4d=52
-4d=52-4
-4d=48
d=48/-4
d=-12 ans.
Proof:
3*4-12=0
12-12=0
0=0